OpenStudy (anonymous):

(Essay Worth 8 points) The formula for degrees Celsius is: C = 9 over 5 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is. ( last one for today please help me answer this qestion ! ?

OpenStudy (anonymous):

Basically, you just have to get to the point where F = something. So, start by looking at the original equation: C = (5/9)*(F-32) Divide both sides by (5/9) Remember, when dividing fractions, switch the numerator and denominator, and then multiply by that fraction. C*(9/5) = F-32. Then add 32: C*(9/5) + 32 = F The 2nd part indicates that you plug a value for C to get the F value: [(6)*9/5] + 32 = F 10.8 + 32 = F 42.8 = F

OpenStudy (anonymous):

heres easier explanation:

OpenStudy (anonymous):

nvm there u go

OpenStudy (anonymous):

thankyou so much :)